# Schwinger dark matter production

### v1

@@ IV Summary and Discussion
We have shown that the observed dark matter relic abundance can be generated via a dark Schwinger mechanism which takes place during inflation and/or after reheating. This mechanism leads to a feebly interacting dark sector comprised of dark fermions and/or scalars which can be as light as $\sim 10\,MeV$ and as heavy as $\sim 10^{10}\,GeV$ depending on how the Schwinger current redshifts after inflation (see Fig. 1). It requires larger Hubble scales at the end of inflation in the range $H_{\rm end}\sim(10^{8}-10^{14})\,GeV$ and reheat temperatures which are not too much lower. The necessary background dark electric field is generated during inflation and can in principle continue sourcing Schwinger production of dark charged particles after reheating until they become non-relativistic. The dark matter can exchange ultralight dark photons and has a power spectrum which is peaked at very small scales, thus evading potential isocurvature constraints from measurements of the CMB.
@@
We have discussed possible sources for the constant background dark electric field during inflation with particular focus on the mechanism introduced in Bastero-Gil et al. (2019, 2022). We have derived the energy density of the Schwinger pair produced dark charged particles at the end of inflation (see Eq. (9)) as well as at anytime after reheating (see Eq. (II)). We then derived the relic abundance of the dark sector today (see Eq. (III)) and examined contours where it matches the observed dark matter relic abundance. We have emphasized that in inflationary Schwinger production the background electric field serves as a source of conformal symmetry breaking which is present even in the absence of conformal breaking in the fermion and scalar sectors. Thus even massless fermions or conformally coupled massless scalars, which would not be produced by gravitational particle production due to a changing metric, can still be produced during inflation (and after) via the inflationary Schwinger effect.
Our analysis relied on analytic calculations of the conductivity associated with the conserved Schwinger current during inflation computed in various studies of inflationary magnetogenesis Hayashinaka et al. (2016); Banyeres et al. (2018). The conductivity after reheating has yet to be computed so in this regime we have considered different possibilities for how the conductivity redshifts. Eventually computing the conductivity after inflation will be necesarry to make robust predictions and an effort in this direction is ongoing.
@@
Phenomenological implications of the non-thermal dark matter sector considered here depend on how the dark sector communicates with the Standard Model, if at all. For instance a coupling to the Standard Model could occur through kinetic mixing with the photon. Another possibility is through Higgs mixing or through the neutrino sector. Even if the dark sector does not communicate with the Standard Model it could have implications for intermediate scale structure or gravitational waves associated with phase transitions in the dark sector. Furthermore, we can have a multi-component dark sector composed of fermions/scalars exchanging an ultralight vector which could have interesting effects on dark matter phenomenology. All of these possibilities are potentially interesting and deserve further investigation.
@@
Acknowledgments: The authors thank Antonio Torres Manso, Manel Masip, Jose Santiago and Takeshi Kobayashi for useful comments and discussions. This work has been partially supported by Junta de Andalucá Project A-FQM-472-UGR20 (fondos FEDER) (R.V.M.) and by SRA (10.13039/501100011033) and ERDF under grant PID2022-139466NB-C21(R.V.M.) as well as by MICINN (PID2019-105943GB-I00/AEI/10.13039/501100011033,PID2022.140831NB.I00 /AEI/10.13039/501100011033/FEDER,UE) (M.B.G.),  FCT-CERN grant No. CERN/FIS-PAR/0027/2021 (M.B.G., P.F.), FCT Grant No.SFRH/BD/151475/2021 (P.F.). The work of LU was supported by the Slovenian Research Agency under the research core funding No. P1-0035 and in part by the research grant J1-4389. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).

### v2

@@ IV Summary and Discussion
We have presented an inflationary Schwinger production mechanism which can generate dark matter during inflation even if purely gravitational production is negligible. We have shown (see Fig. 1 and Fig. 2) that the observed dark matter relic abundance can be generated via a dark Schwinger mechanism based on a ‘dark QED’ $U(1)$ gauge theory which takes place during inflation. This mechanism leads to a feebly interacting dark sector comprised of dark charged fermions and/or scalars which can be as light as $m_{\chi}\sim 0.1$ eV for scalars and $m_{\chi}\sim 100$ keV for fermions and as heavy as $m_{\chi}\sim 10^{12}$ GeV in both cases. In the small mass limit we consider the dark photon which makes up the electric field contributes negligibly to the relic abundance, but must at some point obtain a mass $m_{A}\gtrsim 10^{-10}$ eV in order to evade constraints from structure formation. The dark gauge coupling can have a range $10^{-10}\lesssim g_{D}\lesssim 1$ while the Hubble scale at the end of inflation can span $100\,{\rm GeV}\lesssim H_{\rm end}\lesssim 10^{13}$ GeV with reheating temperatures in a similar range. Overall, we see a very large parameter space for which inflationary Schwinger production can accommodate the observed dark matter relic abundance.
@@
The necessary background dark electric field is generated during inflation and sources Schwinger production until the end of inflation. We have discussed possible sources for this dark electric field during inflation with particular focus on the mechanism introduced in Bastero-Gil et al. (2019, 2022). We have derived the energy density of the Schwinger pair produced dark charged particles at the end of inflation (see Eq. (9)). We then derived the relic abundance of the dark sector today (see Eq. (41)) and examined contours where it matches the observed dark matter relic abundance.
We have emphasized that in inflationary Schwinger production the background electric field serves as a source of conformal symmetry breaking which is present even in the absence of conformal breaking in the fermion and scalar sectors. Thus even massless fermions or conformally coupled massless scalars, which would not be produced by purely gravitational particle production, can still be produced during inflation via the Schwinger effect. The dark matter has a power spectrum which is peaked at very small scales, thus evading potential isocurvature constraints from measurements of the CMB. Our analysis relied on analytic calculations of the conductivity associated with the conserved Schwinger current during inflation computed in various studies of inflationary magnetogenesis Hayashinaka et al. (2016); Banyeres et al. (2018).
@@
The non-thermal dark QED sector contained in this mechanism could have a number of interesting phenomenological possibilities depending on how the dark sector communicates with the Standard Model, if at all. For instance a coupling to the Standard Model could occur through kinetic mixing with the photon. Another possibility is through Higgs mixing or through the neutrino sector. Even if the dark sector does not communicate with the Standard Model it could have implications for intermediate scale structure Archidiacono et al. (2022); Bottaro et al. (2024a, b). Since the dark photon must obtain atleast a small mass there could be gravitational waves associated with a dark Higgs phase transitions in the dark sector. Furthermore, we can have a multi-component dark sector, thermalized with itself or not, composed of fermions/scalars exchanging an ultralight vector which could have interesting effects on dark matter phenomenology. All of these possibilities are potentially interesting and deserve further investigation.
@@
Acknowledgments: The authors thank Antonio Torres Manso, Manel Masip, Jose Santiago and Takeshi Kobayashi for useful comments and discussions. The authors would also like to thank the referees who’s comments led to much improved results and enlarged parameter space. This work has been partially supported by Junta de Andalucá Project A-FQM-472-UGR20 (fondos FEDER) (R.V.M.) and by SRA (10.13039/501100011033) and ERDF under grant PID2022-139466NB-C21(R.V.M.) as well as by MICINN (PID2019-105943GB-I00/AEI/10.13039/501100011033,PID2022.140831NB.I00 /AEI/10.13039/501100011033/FEDER,UE) (M.B.G.),  FCT-CERN grant No. CERN/FIS-PAR/0027/2021 (M.B.G., P.F.), FCT Grant No.SFRH/BD/151475/2021 (P.F.). The work of LU was supported by the Slovenian Research Agency under the research core funding No. P1-0035 and in part by the research grant J1-4389. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). RVM is grateful to the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and partial support during the completion of this work.
V Appendix
In this Appendix we study the dark electric field after inflation and determine how the dark electron momentum evolves. We are interested in the tachyonic production of the dark photon discussed above (see Eq. (12)) in which only one transverse component is produced which we call $A_{+}$. We work with the FLRW metric,
$ds=dt^{2}-a(t)^{2}d\vec{x}^{2}=a(\tau)^{2}(d\tau^{2}-d\vec{x}^{2})\,,$ (44)
where $a(t)$ is the scale factor, and $\tau$ the conformal time, $d\tau=dt/a(t)$. The vector potential can be separated into transverse $\vec{A}_{T}$ and longitudinal $A_{L}$ components, such that $\vec{k}\cdot\vec{A}=kA_{L}$ and $\vec{k}\cdot\vec{A}_{T}=0$. Using the notation $\pm$ for the 2 transverse components, the (classical) vector field can be written as,
$\vec{A}=\sum_{\lambda=\pm,L}\int\frac{d^{3}k}{(2\pi)^{3/2}}(\vec{e}_{\lambda}A_{\lambda}(t)e^{i\vec{k}\cdot\vec{x}}+h.c.)\,,$ (45)
where $\vec{e}_{\lambda}$ are the longitudinal/transverse polarization unit vectors,
$\vec{k}\cdot\vec{e}_{L}=k\,,\;\;\;\vec{k}\cdot\vec{e}_{\pm}=0\,,\;\;\;\vec{k}\times\vec{e}_{\pm}=\mp ik\vec{e}_{\pm}\,.$ (46)
Since the electric and magnetic fields are defined as,
$a^{2}\vec{B}=\vec{\nabla}\times\vec{A}\,,\;\;\;a^{2}\vec{E}=-\partial_{\tau}\vec{A}=-a\partial_{t}\vec{A}\,,$ (47)
the last condition in Eq. (46) ensures that the (transverse) electric and magnetic fields produced during inflation are (anti)parallel. Thus the presence of the magnetic field would only serve to enhance the Schwinger effect Fujita et al. (2022).
Taking for example the comoving momentum in the $z$ direction, $\vec{k}=(0,0,k)$, the longitudinal and transverse unit vectors can be chosen as,
$\vec{e}_{L}$ $=$ $(0,0,1)\,,$ (48)
$\vec{e}_{+}$ $=$ $\frac{1}{\sqrt{2}}(1,i,0)\,,$ (49)
$\vec{e}_{-}$ $=$ $\frac{1}{\sqrt{2}}(i,-1,0)\,.$ (50)
The transverse potential vector components are then given by:
$A_{+}^{x}$ $=$ $\frac{1}{\sqrt{2}}\int_{k}(A_{+}e^{i\vec{k}\cdot\vec{x}}+h.c)$ (51)
$=$ $\frac{2}{\sqrt{2}}Re\int_{k}A_{+}e^{i\mathbf{k}\cdot\vec{x}}=\int_{k}A^{R}_{+}\cos(kz)\,,$
$A_{+}^{y}$ $=$ $\frac{i}{\sqrt{2}}\int_{k}(A_{+}e^{i\vec{k}\cdot\vec{x}}-h.c)=-\int_{k}A^{I}_{+}\sin(kz)\,,$ (52)
and similarly for their time derivatives. From numerical simulations, once the perturbation re-enters the horizon ($k/a>H$) and the field starts oscillating we find that $A_{+}^{R}\sim\pm A_{+}^{I}$, $\dot{A}_{+}^{R}\sim\pm\dot{A}_{+}^{I}$.
The spectrum of $A_{+}$ modes is peaked at the end of inflation around $k_{P}\simeq\mathcal{O}(a_{end}H_{end})$, so let us approximate their amplitudes by a delta function around the peak $k_{P}$ times an oscillatory function with frequency $k\tau$,
$A_{+}^{x}$ $=$ $A_{+}(0)\cos(\Omega_{k})\cos(kz)\,,$ (53)
$A_{+}^{y}$ $=$ ±$\pm A_{+}(0)\cos(\Omega_{k})\sin(kz)\,,$ (54)
$A_{+}(0)$ being the initial amplitude of the oscillations and,
$\Omega_{k}\simeq k_{P}\tau=k_{P}\int\frac{dt}{a}=k_{P}\int\frac{da}{a^{2}H}=\frac{k_{P}}{H_{0}}(\frac{a}{a_{0}}-1)\,,~{}~{}~{}~{}$ (55)
where in the last equality we have assumed that the universe is radiation dominated. The physical electric field, given by $\vec{E}=-\partial_{\tau}\vec{A}=-\partial_{t}\vec{A}/a$, would be given by,
$E^{x}\simeq E^{x}_{+}(t)\sin(\Omega_{k}(t))\cos(kz)\,,$ (56)
$E^{y}\simeq\pm E^{y}_{+}(t)\sin(\Omega_{k}(t))\sin(kz)\,,$ (57)
$E^{i}_{+}$ being the amplitude of the oscillations. We have kept the superindex $i=x,\,y$ because the initial amplitudes may not be exactly the same. We know that once inflation and the tachyonic resonance end, the energy density of the dark vector will behave as radiation and therefore [^5] We use the subindex “0” to denote the values at the start of the oscillations. $E^{i}_{+}=E^{i}_{0}/a^{2}$.
The physical momentum of the charged particles follows the evolution equation,
$\frac{dp^{i}}{dt}+Hp^{i}\simeq g_{D}E^{i}\,,$ (58)
where we have neglected the (subdominant) magnetic field contribution. For example, for the “x” component, and dividing by $H$, we have,
$\frac{dp^{x}}{d\ln a}+p^{x}=\frac{g_{D}E^{x}_{0}}{H_{0}}\sin(\Omega_{k}(a/a_{0}))\cos(kz)\,,$ (59)
where again we have used the fact that both $E_{+}(t)$ and $H(t)$ redshift as $a^{-2}$. Using $p^{x}=\bar{p}^{x}/a$ and $u=a/a_{0}$ we have:
$\frac{d\bar{p}^{x}}{du}=\frac{g_{D}E^{x}_{0}}{H_{0}}\sin(\frac{k_{P}}{H_{0}}(u-1))\cos(kz)\,,$ (60)
and then,
$p^{x}$ $=$ $\frac{a_{0}}{a}\bar{p}^{x}$ (61)
$=$ $\frac{a_{0}}{a}\left[p^{x}(0)+\frac{g_{D}E^{x}_{0}}{k_{P}}2\sin^{2}(\frac{k_{P}}{H_{0}}(\frac{a}{a_{0}}-1))\cos(kz)\right]\,,~{}~{}~{}~{}~{}$
where $p^{x}(0)$ is the initial value. Similarly for the other component,
$p^{y}$ $=$ $\frac{a_{0}}{a}\bar{p}^{y}$ (62)
$=$ $\frac{a_{0}}{a}\left[p^{y}(0)\pm\frac{g_{D}E^{y}_{0}}{k_{P}}2\sin^{2}(\frac{k_{P}}{H_{0}}(\frac{a}{a_{0}}-1))\sin(kz)\right]\,.~{}~{}~{}~{}~{}$
Therefore, the physical momentum is always redshifted with the expansion.
As an example in Fig. 3 we show the time evolution after inflation of a couple of modes (the real and imaginary components) of the vector potential and its time derivative. We have used a quartic chaotic potential $V=\lambda\phi/4$, with $\lambda=10^{-14}$, and $\alpha m_{P}/f=7.5$. The integration starts before inflation ends, but we only show the evolution from $\epsilon_{H}=1$ ($N_{e}=0$) onwards. In the plot it is indicated the value of the comoving momentum in units of the $a_{e}H_{e}$, where $H_{e}$ is the value of the Hubble parameter at the end of inflation.
Figure 3: